OpenStudy (anonymous):

Does anyone know how to solve Base 10 logarithm problems?

OpenStudy (vishal_kothari):

what problem r u facing in solving it?

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

18^(x-2) = 13^(-2x) Taking logarithm to the base 10, (x-2) log 18 = -2x log 13 or x log 18 + 2x log 13 = 2 log 18 or (log 18 + 2 log 13) x = 2 log 18 or x = (2 log 18)/(log 18 + 2 log 13)

OpenStudy (anonymous):

Probably

OpenStudy (unklerhaukus):

\[~~~~~~~~~~~~~18^{(x-2)} = 13^{(-2x)}\]\[~~~~~log_{10}18^{(x-2)} = log_{10}13^{(-2x)}\]\[(x-2)log_{10}18 = (-2x)log_{10}13\]\[(x-2)log_{10}18 + (2x)log_{10}13=0\]\[xlog_{10}18 + 2xlog_{10}13=2log_{10}18 \]\[x(log_{10}18+2log_{10}13)=2log_{10}18 \]\[x={2log_{10}18 \over log_{10}18+2log_{10}13}\] all are correct